OpenStudy (thedj4jc):

Reduce the fraction to lowest terms. Do not use spaces in your answer. -5jk ------- 35j^2k^2

OpenStudy (anonymous):

start by simplifying the -5/35

OpenStudy (shamim):

-5/5=?

OpenStudy (shamim):

35/5=?

OpenStudy (thedj4jc):

-1 7

OpenStudy (anonymous):

then you need to look at the variables and remove any extras

OpenStudy (anonymous):

i got the answer

OpenStudy (shamim):

So we can write -5/35=-1/7 Right???

OpenStudy (anonymous):

want me to tell you.......?

OpenStudy (anonymous):

@shamim is wrong

OpenStudy (shamim):

I did not finish solution

OpenStudy (thedj4jc):

@shamim Yes, we can.

OpenStudy (shamim):

j/j^2=?

OpenStudy (anonymous):

the correct answer is \[-\frac{ 1 }{ 7jk}\]

OpenStudy (anonymous):

the variables are the easy part. What is j/j^2?

OpenStudy (thedj4jc):

That'd be just j, right?

OpenStudy (anonymous):

it would by 1/j

OpenStudy (anonymous):

my answer is correct

OpenStudy (shamim):

Ya j/j^2=1/j

OpenStudy (anonymous):

@pjpineda17 is correct but shes supposed to show work lol

OpenStudy (shamim):

Now k/k^2=?

OpenStudy (thedj4jc):

1/k

OpenStudy (shamim):

Right u r!!!!!!!

OpenStudy (anonymous):

hehe sorry i used a website called https://mathway.com its very helpful

OpenStudy (shamim):

So the result is after simplification -1/(7jk) Right????!!!

OpenStudy (thedj4jc):

Ok, thanks y'all!!!

OpenStudy (shamim):

U r most welcome!!!!!!

OpenStudy (anonymous):

can i get a medal